OpenStudy (anonymous):

Josh believes the Spanish club students at his school have an unfair advantage in being assigned to the Spanish class they request. He asked 500 students at his school the following questions: "Are you in the Spanish club?" and "Did you get the Spanish class you requested?" The results are shown in the table below: Spanish Club Not in Spanish Club Total Received Spanish class requested 265 100 365 Did not get Spanish class requested 70 65 135 Total 335 165 500

OpenStudy (anonymous):

Josh believes the Spanish club students at his school have an unfair advantage in being assigned to the Spanish class they request. He asked 500 students at his school the following questions: "Are you in the Spanish club?" and "Did you get the Spanish class you requested?" The results are shown in the table below: Spanish Club Not in Spanish Club Total Received Spanish class requested 265 100 365 Did not get Spanish class requested 70 65 135 Total 335 165 500 Help Josh determine if all students at his school have an equal opportunity to get the Spanish class they requested. Show your work and explain your process for determining the fairness of the class assignment process.

OpenStudy (biue):

Hm

OpenStudy (biue):

P(students in spanish club who recieved spanish class) = 265/365 = 53/73 P(students NOT in spanish club who recieved spanish class) = 100/365 = 20/73 so all students dont have equal opportunity to get the Spanish class they requested

OpenStudy (biue):

the question is saying that we have to find that all the students of Josh's class wether they r in club or not can attend spanish class or not. So we just take the data of people who can attend the spanish class and take out the probilities of that. for finding the probabilty we divide the no. of students of each group who can attend the class by the total no. of students who atteded..

OpenStudy (biue):

if both group probabilities are same then all students have equal opportunity to get the Spanish class otherwise they dont have equal opportunity

OpenStudy (anonymous):

Wait! Continue!

OpenStudy (anonymous):

@satellite73